Site Conditions:
  • No smoke detected on site
  • Odor detected throughout the site
Activities:
  • Material is being moved from the northern area, Section B, to the middle of the site
  • The majority of the southern section, Section A, has been covered with hay and grass seed, and grass is starting to sprout in a few areas
  • Work being performed today and for the next several days will include grading and preparing the rest of the site to be covered with hay and grass seed
  • EPA personnel are collecting air samples for testing at 10 points near the site